# A mathematical theory of topological invariants of quantum spin systems

### v1

@@ 4.1 Semilinear sets and their thickenings
A semilinear set in ${\mathbb{R}}^{n}$ is a subset of ${\mathbb{R}}^{n}$ which can be defined by a finite number of linear equalities and strict linear inequalities [20]. More precisely, a basic semilinear set in ${\mathbb{R}}^{n}$ is an intersection of a finite number of hyperplanes and open half-spaces, and a semilinear set is a finite union of basic semilinear sets [20]. The set of semilinear subsets of ${\mathbb{R}}^{n}$ will be denoted ${\mathcal{S}}_{n}$. Projections ${\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ map ${\mathcal{S}}_{m+n}$ to ${\mathcal{S}}_{m}$ [20].
A map ${\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ is called semilinear iff its graph is a semilinear subset of ${\mathbb{R}}^{m+n}$. A composition of two semilinear maps is a semilinear map [20].
@@
The function $d:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is semilinear.
@@
The function $({\mathbf{x}},{\mathbf{y}})\mapsto x_{i}-y_{i}$ is semilinear for any $i$. The function $|\cdot|:{\mathbb{R}}\rightarrow{\mathbb{R}}$ is semilinear. If $f,g:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ are semilinear, then $h=\max(f,g):{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is semilinear. Since the set of semilinear functions is closed under composition, this proves the lemma. ∎
@@
$\Delta_{r}=\{({\mathbf{x}},{\mathbf{y}})\in{\mathbb{R}}^{2n}|d({\mathbf{x}},{\mathbf{y}})\leq r\}.$
@@
Suppose ${\mathbf{x}},{\mathbf{y}}\in U^{r}$, and suppose ${\mathbf{x}}^{\prime},{\mathbf{y}}^{\prime}\in U$ satisfy $d({\mathbf{x}},{\mathbf{x}}^{\prime})\leq r$ and $d({\mathbf{y}},{\mathbf{y}}^{\prime})\leq r$. Then for any $t\in[0,1]$ we have $d(t{\mathbf{x}}+(1-t){\mathbf{y}},t{\mathbf{x}}^{\prime}+(1-t){\mathbf{x}}^{\prime})\leq t^{2}d({\mathbf{x}},{\mathbf{x}}^{\prime})+(1-t)^{2}d({\mathbf{y}},{\mathbf{y}}^{\prime})\leq r.$
@@ 4.2 A category of fuzzy semilinear sets
We need the following geometric result [22] [^11] The proof in [22] is for the Euclidean distance, but since the Euclidean distance function and $d({\mathbf{x}},{\mathbf{y}})=\|{\mathbf{x}}-{\mathbf{y}}\|_{\infty}$ are equivalent (each one is upper-bounded by a multiple of the other), the result applies to $d({\mathbf{x}},{\mathbf{y}})$ as well. :
@@
$d({\mathbf{x}},U^{r}\cap V^{r})\leq C\max(d({\mathbf{x}},U^{r}),d({\mathbf{x}},V^{r}))$ (85)
for each pair $i,j$. Since $U^{r}=\cup_{i}P_{i}^{r}$ and $V^{r}=\cup_{i}Q_{i}^{r}$, for any ${\mathbf{x}}\in{\mathbb{R}}$ there are indices $i^{*}$ and $j^{*}$ such that $d({\mathbf{x}},U^{r})=d({\mathbf{x}},P_{i^{*}}^{r})$ and $d({\mathbf{x}},V^{r})=d({\mathbf{x}},Q_{j^{*}}^{r})$. Then we have
$d({\mathbf{x}},U^{r}\cap V^{r})$ $=d({\mathbf{x}},\cup_{ij}(P_{i}^{r}\cap Q_{j}^{r}))$ (86)
$\leq d({\mathbf{x}},P_{i^{*}}^{r}\cap Q_{j^{*}}^{r}))$ (87)
$\leq C_{P_{i^{*}}Q_{j^{*}}}\max(d({\mathbf{x}},P_{i^{*}}),d({\mathbf{x}},Q_{j^{*}}))$ (88)
$=C_{P_{i^{*}}Q_{j^{*}}}\max(d({\mathbf{x}},U),d({\mathbf{x}},V)),$ (89)
@@
Since $U$ and $U^{r}$ (resp. $V$ and $V^{r}$) are isomorphic in ${\mathcal{C}\mathcal{S}}_{n}$, it suffices to show that $U^{r}\cap V^{r}$ is a meet of $U^{r}$ and $V^{r}$. It’s clear that $U^{r}\cap V^{r}\leq U^{r}$ and $U^{r}\cap V^{r}\leq V^{r}$. Now suppose $W\in{\mathcal{C}\mathcal{S}}_{n}$ satisfies $W\leq U^{r}$ and $W\leq V^{r}$. Then there is an $s>0$ such that every ${\mathbf{x}}\in W$ satisfies $\max(d({\mathbf{x}},U^{r}),d({\mathbf{x}},V^{r}))\leq s$. Since $d({\mathbf{x}},U^{r}\cap V^{r})\leq C\max(d({\mathbf{x}},U^{r}),d({\mathbf{x}},V^{r}))$ for some $C>0$ we have $W\subset(U^{r}\cap V^{r})^{Cs}$. ∎
@@
We can now equip ${\mathcal{C}\mathcal{S}}_{n}$ with a Grothendieck topology of Section 2. There are two versions of it which differ in whether we allow empty covers of smallest elements or not. In the case of the pre-ordered set ${\mathcal{C}\mathcal{S}}_{n}$, every bounded closed semilinear set is a smallest element (they are all isomorphic objects of the category ${\mathcal{C}\mathcal{S}}_{n}$). Since such sets are not empty, the more natural choice is not to allow empty covers. This choice is also forced on us if we want certain pre-cosheaves to be cosheaves (see below). Note that we only consider non-empty closed semilinear sets.
@@ 4.3 Spherical CS sets
Every two bounded elements of ${\mathcal{C}\mathcal{S}}_{n}$ are isomorphic objects of the corresponding category. More generally, any two elements of ${\mathcal{C}\mathcal{S}}_{n}$ which coincide outside some ball in ${\mathbb{R}}^{n}$ are isomorphic objects. Thus ${\mathcal{C}\mathcal{S}}_{n}$ encodes the large-scale structure of ${\mathbb{R}}^{n}$. To make this explicit, we will show that the pre-ordered set ${\mathcal{C}\mathcal{S}}_{n}$ is equivalent as a category to a certain poset of subsets of the ”sphere at infinity” $S^{n-1}$.
A cone in ${\mathbb{R}}^{n}$ is a non-empty subset of ${\mathbb{R}}^{n}$ which is invariant under ${\mathbf{x}}\mapsto\lambda{\mathbf{x}}$, where $\lambda\geq 0$. Every cone contains the origin 00. Cones in ${\mathbb{R}}^{n}$ are in bijection with subsets of $S^{n-1}=({\mathbb{R}}^{n}\backslash\{0\})/{\mathbb{R}}^{*}_{+}$ where ${\mathbb{R}}^{*}_{+}$ is the group of positive real numbers under multiplication. If $A\subset S^{n-1}$, we denote the corresponding cone $c(A)$. In particular, $c(\varnothing)=\{0\}\in{\mathbb{R}}^{n}$. If $K$ is a cone in ${\mathbb{R}}^{n}$, we will denote the corresponding subset of $S^{n-1}$ by $\hat{K}$.
@@
We proceed first by defining a functor from ${\mathcal{C}\mathcal{S}}_{n}$ to ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. Notice, it is sufficient to define a functor on isomorphism classes of polyhedrons then extend by joins. Every closed polyhedron is a finite intersection of closed half spaces $\bigcap_{i}\{{\mathbf{n}}_{i}\cdot{\mathbf{x}}\leq b\},$ where ${\mathbf{n}}_{i}$’s are unit normal vectors of supporting hyperplanes. By definition, $\bigcap_{i}\{{\mathbf{n}}_{i}\cdot{\mathbf{x}}\leq b\}$ is isomorphic to $\bigcap_{i}\{{\mathbf{n}}_{i}\cdot{\mathbf{x}}\leq 0\}$ in ${\mathcal{C}\mathcal{S}}_{n}$. In other words, every polyhedron is isomorphic to a cone $\{{\mathbf{n}}{\mathbf{x}}\leq 0\},$ where ${\mathbf{n}}$ is the matrix with ${\mathbf{n}}_{i}$ as rows. If $\{{\mathbf{n}}_{i}\}_{i}$ spans ${\mathbb{R}}^{n}$, the set $\{{\mathbf{n}}\ {\mathbf{x}}\leq 0\}$, which contains no antipodal points, is mapped to its corresponding spherical polyhedron. Otherwise, complete $\{{\mathbf{n}}_{i}\}_{i}$ into a spanning set $\{{\mathbf{n}}_{1},\dots,{\mathbf{n}}_{k},{\mathbf{m}}_{1},\dots,{\mathbf{m}}_{l}\}$ and decompose $\{{\mathbf{n}}\ {\mathbf{x}}\leq 0\}$ into a union of cones $\{{\mathbf{n}}\ {\mathbf{x}}\leq 0\}\cap\bigcap_{j}\{\pm{\mathbf{m}}_{j}\cdot{\mathbf{x}}\leq 0\}$. By the universal property of the join, an arbitrary closed semilinear set is isomorphic to a union of conical polyhedra. For functoriality, it is sufficient to note that for any pair of closed cones $U,V$ a fuzzy inclusion $U\subseteq V^{r}$ for some $r\geq 0$ implies $U\subseteq V$. It is easy to check that this functor is fully faithful and (essentially) surjective. ∎
@@ 5.1 Basic examples
By Lemma 3.2 the maps sending $U\in{\mathcal{C}\mathcal{S}}_{n}$ to ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}^{\psi}(U)$ for a given state $\psi$ are pre-cosheaves of Fréchet spaces on ${\mathcal{C}\mathcal{S}}_{n}$. Denote these pre-cosheaves by ${\mathfrak{D}}_{al}$, and ${\mathfrak{D}}_{al}^{\psi}$. Putting together Lemma 5.1, Proposition 3.5, Corollary 3.1, and Proposition 3.8, we have
@@
${\mathfrak{D}}_{al}$ is a real local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$. If $\psi$ is gapped, then ${\mathfrak{D}}_{al}^{\psi}$ is a real local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$.
@@
$p_{k,U}(f)=\sup_{j\in\Lambda}|f(j)|(1+d(U,j))^{k},\quad k\in{\mathbb{N}}.$ (90)
It is easy to check that the assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ it is a coflasque pre-cosheaf of Fréchet-Lie algebras satisfying Property I. To check the cosheaf property, the following lemma is useful.
Lemma 5.1.
Let ${\mathfrak{F}}$ be a pre-cosheaf of vector spaces on a distributive lattice $L$, with co-restriction maps $\iota_{U,V}:U\to V$. Then ${\mathfrak{F}}$ is a cosheaf on $L$ (with the topology of finite covers) iff for any $U,V\in L$ the following sequence is exact
[…]
@@
We must show that for any $U\in L$ and every covering $U_{1}\vee\ldots\vee U_{n}=U$, the sequence of vector spaces
[…]
is exact, where $\alpha=\sum_{i<j}\iota_{U_{i}\wedge U_{j},U_{i}}-\iota_{U_{i}\wedge U_{j},U_{j}}$ and $\beta=\sum_{i}\iota_{U_{i},U}$. Exactness of the last three terms follows from an easy induction and the associativity of the join operation. For exactness of the first three terms of the sequence, we also proceed by induction: suppose the result holds for all covers of cardinality $n-1$, and let $U_{1},\ldots,U_{n}\in{\mathcal{C}\mathcal{S}}_{n}$ be a cover of $U$. Suppose $(s_{1},\ldots,s_{n})\in\bigoplus_{i=1}^{n}{\mathfrak{F}}(U_{i})$ lies in the kernel of $\beta$. Let $V:=\bigvee_{i=1}^{n-1}U_{i}$. An application of (91) with $U_{n}$ and $V$ shows that $s_{n}=\iota_{U_{n}\cap V,U}(t)$ for some $t\in V\wedge U_{n}=(U_{1}\wedge U_{n})\vee...\vee(U_{n-1}\vee U_{n})$, and by right-exactness of (92) this shows that $s_{n}=\sum_{i=1}^{n-1}\iota_{U_{i}\cap U_{n},U_{n}}(v_{i})$ for some $v_{i}\in{\mathfrak{F}}(U_{i})$. Finally we write
$(s_{1},\ldots,s_{n})=(s_{1}+w_{1},\ldots,s_{n-1}+w_{n-1},0)+(-w_{1},\ldots,-w_{n-1},s_{n})$ (93)
Proposition 5.1.
The assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ is a cosheaf of vector spaces.
Proof.
By the preceding lemma, it is sufficient to show that for any $U,V\in{\mathcal{C}\mathcal{S}}_{n}$ the sequence
${{\mathfrak{g}}_{al}}(U\wedge V)\rightarrow{{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)\rightarrow{{\mathfrak{g}}_{al}}(U\vee V)\rightarrow 0$ (94)
@@
$f_{U}({\mathbf{x}})=\begin{cases}f({\mathbf{x}}),&d({\mathbf{x}},U)<d({\mathbf{x}},V),\\ \frac{1}{2}f({\mathbf{x}}),&d({\mathbf{x}},U)=d({\mathbf{x}},V),\\ 0,&d({\mathbf{x}},U)>d({\mathbf{x}},V),\end{cases}$ (95)
and $f_{V}({\mathbf{x}})$ is defined by a similar expression with $U$ and $V$ exchanged. Using $d({\mathbf{x}},U\cup V)=\min(d({\mathbf{x}},U),d({\mathbf{x}},V))$ it is easy to check that $p_{k,U}(f_{U})\leq p_{k,U\cup V}(f)$ and $p_{k,V}(f_{V})\leq p_{k,U\cup V}(f)$ for all $k$, and thus $f_{U}\in{{\mathfrak{g}}_{al}}(U)$ and $f_{V}\in{{\mathfrak{g}}_{al}}(V)$.
To show exactness at ${{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)$, suppose $f\in{{\mathfrak{g}}_{al}}(U)\cap{{\mathfrak{g}}_{al}}(V)$. From the proof of Prop. 4.2, there exist $r\geq 0,C_{UV}>0$ such that $d({\mathbf{x}},U^{r}\cap V^{r})\leq C_{UV}\max(d({\mathbf{x}},U),d({\mathbf{x}},V))$. We may assume that $C_{UV}\geq 1$, in which case for any ${\mathbf{x}}\in{\mathbb{R}}^{n}$ and any $k\in{\mathbb{N}}$
$(1+d({\mathbf{x}},U^{r}\cap V^{r}))^{k}\leq C^{k}_{UV}(1+\max(d({\mathbf{x}},U),d({\mathbf{x}},V))^{k}.$ (96)

### v2

@@ 4.1 Semilinear sets and their thickenings
A semilinear set in ${\mathbb{R}}^{n}$ is a subset of ${\mathbb{R}}^{n}$ which can be defined by means of a finite number of linear equalities and strict linear inequalities. More precisely, a basic semilinear set in ${\mathbb{R}}^{n}$ is an intersection of a finite number of hyperplanes and open half-spaces, and a semilinear set is a finite union of basic semilinear sets. The set of semilinear subsets of ${\mathbb{R}}^{n}$ will be denoted ${\mathcal{S}}_{n}$. Projections ${\mathbb{R}}^{m}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ map ${\mathcal{S}}_{m+n}$ to ${\mathcal{S}}_{m}$ [23]. A map ${\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{m}$ is called semilinear iff its graph is a semilinear subset of ${\mathbb{R}}^{m+n}$. The composition of two semilinear maps is a semilinear map [23].
@@
The distance function $d:{\mathbb{R}}^{n}\times{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is semilinear.
@@
The function $(x,y)\mapsto x_{i}-y_{i}$ is semilinear for any $i$. The function $|\cdot|:{\mathbb{R}}\rightarrow{\mathbb{R}}$ is semilinear. If $f,g:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ are semilinear, then $h=\max(f,g):{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is semilinear. Since the set of semilinear functions is closed under composition, this proves the lemma. ∎
@@
$\Delta_{r}=\{(x,y)\in{\mathbb{R}}^{2n}|d(x,y)\leq r\}.$
@@
Suppose $x,y\in U^{r}$, and suppose $x^{\prime},y^{\prime}\in U$ satisfy $d(x,x^{\prime})\leq r$ and $d(y,y^{\prime})\leq r$. Then for any $t\in[0,1]$ we have $d(tx+(1-t)y,tx^{\prime}+(1-t)x^{\prime})\leq t^{2}d(x,x^{\prime})+(1-t)^{2}d(y,y^{\prime})\leq r.$
@@ 4.2 A category of fuzzy semilinear sets
We need the following geometric result [25] [^9] The proof in [25] is for the Euclidean distance, but since the Euclidean distance function and $d(x,y)=\|x-y\|_{\infty}$ are equivalent (each one is upper-bounded by a multiple of the other), the result applies to $d(x,y)$ as well. :
@@
$d(x,U^{r}\cap V^{r})\leq C\max(d(x,U^{r}),d(x,V^{r}))$
for each pair $i,j$. Since $U^{r}=\cup_{i}P_{i}^{r}$ and $V^{r}=\cup_{i}Q_{i}^{r}$, for any $x\in{\mathbb{R}}$ there are indices $i^{*}$ and $j^{*}$ such that $d(x,U^{r})=d(x,P_{i^{*}}^{r})$ and $d(x,V^{r})=d(x,Q_{j^{*}}^{r})$. Then we have
$d(x,U^{r}\cap V^{r})$ $=d(x,\cup_{ij}(P_{i}^{r}\cap Q_{j}^{r}))$
$\leq d(x,P_{i^{*}}^{r}\cap Q_{j^{*}}^{r}))$
$\leq C_{P_{i^{*}}Q_{j^{*}}}\max(d(x,P_{i^{*}}),d(x,Q_{j^{*}}))$
$=C_{P_{i^{*}}Q_{j^{*}}}\max(d(x,U),d(x,V)),$
@@
Since $U$ and $U^{r}$ (resp. $V$ and $V^{r}$) are isomorphic in ${\mathcal{C}\mathcal{S}}_{n}$, it suffices to show that $U^{r}\cap V^{r}$ is a meet of $U^{r}$ and $V^{r}$. It’s clear that $U^{r}\cap V^{r}\leq U^{r}$ and $U^{r}\cap V^{r}\leq V^{r}$. Now suppose $W\in{\mathcal{C}\mathcal{S}}_{n}$ satisfies $W\leq U^{r}$ and $W\leq V^{r}$. Then there is an $s>0$ such that every $x\in W$ satisfies $\max(d(x,U^{r}),d(x,V^{r}))\leq s$. Since $d(x,U^{r}\cap V^{r})\leq C\max(d(x,U^{r}),d(x,V^{r}))$ for some $C>0$ we have $W\subset(U^{r}\cap V^{r})^{Cs}$. ∎
@@
We can now equip ${\mathcal{C}\mathcal{S}}_{n}$ with a Grothendieck topology of Section 2. There are two versions of it which differ in whether we allow empty covers of an initial object or not. In the case of the pre-ordered set ${\mathcal{C}\mathcal{S}}_{n}$, every bounded closed semilinear set is an initial object (they are all isomorphic objects of the category ${\mathcal{C}\mathcal{S}}_{n}$). Since such sets are not empty, we will disallow empty covers. This choice is also forced on us if we want certain pre-cosheaves to be cosheaves (see below). Note that we only consider non-empty closed semilinear sets.
@@ 4.3 Spherical CS sets
Every two bounded elements of ${\mathcal{C}\mathcal{S}}_{n}$ are isomorphic objects of the corresponding category. More generally, any two elements of ${\mathcal{C}\mathcal{S}}_{n}$ which coincide outside some ball in ${\mathbb{R}}^{n}$ are isomorphic objects. Thus ${\mathcal{C}\mathcal{S}}_{n}$ encodes the large-scale structure of ${\mathbb{R}}^{n}$. To make this explicit, we will show that the pre-ordered set ${\mathcal{C}\mathcal{S}}_{n}$ is equivalent as a category to a certain poset of subsets of the “sphere at infinity” $S^{n-1}$.
A cone in ${\mathbb{R}}^{n}$ is a non-empty subset of ${\mathbb{R}}^{n}$ which is invariant under $x\mapsto\lambda x$, where $\lambda\geq 0$. Every cone contains the origin 00. Cones in ${\mathbb{R}}^{n}$ are in bijection with subsets of $S^{n-1}=({\mathbb{R}}^{n}\backslash\{0\})/{\mathbb{R}}^{*}_{+}$ where ${\mathbb{R}}^{*}_{+}$ is the group of positive real numbers under multiplication. If $A\subset S^{n-1}$, we denote the corresponding cone $c(A)$. In particular, $c(\varnothing)=\{0\}\in{\mathbb{R}}^{n}$. If $K$ is a cone in ${\mathbb{R}}^{n}$, we will denote the corresponding subset of $S^{n-1}$ by $\hat{K}$.
@@
We proceed first by defining a functor from ${\mathcal{C}\mathcal{S}}_{n}$ to ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. Notice, it is sufficient to define a functor on isomorphism classes of polyhedrons then extend by joins. Every closed polyhedron is a finite intersection of closed half spaces $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq b_{i}\},$ where for each $i=1,\ldots k$, $n_{i}\in{\mathbb{R}}^{n}$ is the unit normal vector of the $i$th supporting hyperplane and $b_{i}\in{\mathbb{R}}$. By definition, $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq b_{i}\}$ is isomorphic to $\bigcap_{i=1}^{k}\{n_{i}\cdot x\leq 0\}$ in ${\mathcal{C}\mathcal{S}}_{n}$. In other words, every polyhedron is isomorphic to a cone $\{Nx\preceq 0\},$ where $N$ is the matrix with $n_{i}$ as rows and we write $y\preceq 0$ for $y\in{\mathbb{R}}^{n}$ when $y_{i}\leq 0$ for all $i=1,\ldots n$. If $\{n_{i}\}_{i=1}^{k}$ spans ${\mathbb{R}}^{n}$, the set $\{Nx\preceq 0\}$, which contains no antipodal points, is mapped to its corresponding spherical polyhedron. Otherwise, complete $\{n_{i}\}_{i=1}^{k}$ into a spanning set $\{n_{1},\dots,n_{k},m_{1},\dots,m_{l}\}$ and decompose $\{Nx\preceq 0\}$ into a union of cones $\{Nx\preceq 0\}\cap\bigcap_{j}\{\pm m_{j}\cdot x\leq 0\}$. By the universal property of the join, an arbitrary closed semilinear set is isomorphic to a union of conical polyhedra. For functoriality, it is sufficient to note that for any pair of closed cones $U,V$ a fuzzy inclusion $U\subseteq V^{r}$ for some $r\geq 0$ implies $U\subseteq V$. It is easy to check that this functor is fully faithful and (essentially) surjective. ∎
@@ 5.1 Basic examples
Let $\psi$ be a state of a quantum lattice system on ${\mathbb{R}}^{n}$. By Lemma 3.2 the maps sending $U\in{\mathcal{C}\mathcal{S}}_{n}$ to ${\mathfrak{D}}_{al}(U)$ and ${\mathfrak{D}}_{al}^{\psi}(U)$ are pre-cosheaves of Fréchet spaces on ${\mathcal{C}\mathcal{S}}_{n}$. We denote these pre-cosheaves by ${\mathfrak{D}}_{al}$ and ${\mathfrak{D}}_{al}^{\psi}$. Putting together Lemma 2.1, Proposition 3.5, Corollary 3.1, and Proposition 3.8, we have
@@
${\mathfrak{D}}_{al}$ is a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$. If $\psi$ is gapped, then ${\mathfrak{D}}_{al}^{\psi}$ is a local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$.
@@
$p_{k,U}(f)=\sup_{j\in\Lambda}|f(j)|(1+d(U,j))^{k},\quad k\in{\mathbb{N}}.$
Proposition 5.1.
The assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ is a local Lie algebra.
@@
It is easy to check that the assignment $U\mapsto{{\mathfrak{g}}_{al}}(U)$ it is a coflasque pre-cosheaf of Fréchet-Lie algebras satisfying Property I. The only thing left to check is that it is a cosheaf of vector spaces. By Lemma 2.1, it is sufficient to show that for any $U,V\in{\mathcal{C}\mathcal{S}}_{n}$ the sequence
${{\mathfrak{g}}_{al}}(U\wedge V)\rightarrow{{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)\rightarrow{{\mathfrak{g}}_{al}}(U\vee V)\rightarrow 0$
@@
$f_{U}(x)=\begin{cases}f(x),&d(x,U)<d(x,V),\\ \frac{1}{2}f(x),&d(x,U)=d(x,V),\\ 0,&d(x,U)>d(x,V),\end{cases}$
and $f_{V}(x)$ is defined by a similar expression with $U$ and $V$ exchanged. Using $d(x,U\cup V)=\min(d(x,U),d(x,V))$ it is easy to check that $p_{k,U}(f_{U})\leq p_{k,U\cup V}(f)$ and $p_{k,V}(f_{V})\leq p_{k,U\cup V}(f)$ for all $k$, and thus $f_{U}\in{{\mathfrak{g}}_{al}}(U)$ and $f_{V}\in{{\mathfrak{g}}_{al}}(V)$.
To show exactness at ${{\mathfrak{g}}_{al}}(U)\oplus{{\mathfrak{g}}_{al}}(V)$, suppose $f\in{{\mathfrak{g}}_{al}}(U)\cap{{\mathfrak{g}}_{al}}(V)$. From the proof of Prop. 4.2, there exist $r\geq 0,C_{UV}>0$ such that $d(x,U^{r}\cap V^{r})\leq C_{UV}\max(d(x,U),d(x,V))$. We may assume that $C_{UV}\geq 1$, in which case for any $x\in{\mathbb{R}}^{n}$ and any $k\in{\mathbb{N}}$
$(1+d(x,U^{r}\cap V^{r}))^{k}\leq C^{k}_{UV}(1+\max(d(x,U),d(x,V))^{k}.$